OpenStudy (anonymous):

Stumped one this one. A piece of cake for you guys, probably haha so yeah here it is. Solve the linear system using substitution. 3x+2y=5 5x-9y=-4 Thanks. :)

OpenStudy (harsimran_hs4):

find out y in terms of x or x in terms of y and substitute it`s value in the 2nd equation

OpenStudy (anonymous):

Yep, and I got y=3/2x + 5/2 Plugged it into the second equation but came out with a totally wack answer.

OpenStudy (harsimran_hs4):

1st part is wrong it should be y = -3/2x +5/2 you missed the - sign

OpenStudy (anonymous):

Nooo, pretty sure there isn't negative in the first one. :P

OpenStudy (harsimran_hs4):

3x+2y=5 2y = 5 - 3x y = 5/2 - 3/2x so it`s - sign whereas you have written y = 5/2 +3/2x which is wrong

OpenStudy (anonymous):

Oh yeah, we'll I did do a negative on paper, re wrote it wrong here.

OpenStudy (anonymous):

You get... 5x-9(-3/2x+5/2) = -4

OpenStudy (harsimran_hs4):

yes this is right

OpenStudy (anonymous):

Then I got.... 5x-27/2x-45/2=-4

OpenStudy (harsimran_hs4):

no - with - cancels out so it should be 5x + 27/2x - 45/2=-4

OpenStudy (anonymous):

Indeed, sorry. Then I got, 37/2x = 18 1/2 I may lose you at this part, so yeah brace yourself. :P

OpenStudy (harsimran_hs4):

yes thats right and it gives you x = 1

OpenStudy (anonymous):

That's correct, thanks! :-D

OpenStudy (harsimran_hs4):

welcome :)